# Patho-R1: A Multimodal Reinforcement Learning-Based Pathology Expert Reasoner

### v1

@@ Abstract
Recent advances in vision-language models (VLMs) have enabled broad progress in the general medical field. However, pathology still remains a more challenging sub-domain, with current pathology-specific VLMs exhibiting limitations in both diagnostic accuracy and reasoning plausibility. Such shortcomings are largely attributable to the nature of current pathology datasets, which are primarily composed of image–description pairs that lack the depth and structured diagnostic paradigms employed by real-world pathologists. In this study, we leverage pathology textbooks and real-world pathology experts to construct high-quality, reasoning-oriented datasets. Building on this, we introduce Patho-R1, a multimodal RL-based pathology Reasoner, trained through a three-stage pipeline: (1) continued pretraining on 3.5 million image-text pairs for knowledge infusion; (2) supervised fine-tuning on 500k high-quality Chain-of-Thought samples for reasoning incentivizing; (3) reinforcement learning using Group Relative Policy Optimization and Decoupled Clip and Dynamic sAmpling Policy Optimization strategies for multimodal reasoning quality refinement. To further assess the alignment quality of our dataset, we propose PathoCLIP, trained on the same figure-caption corpus used for continued pretraining. Comprehensive experimental results demonstrate that both PathoCLIP and Patho-R1 achieve robust performance across a wide range of pathology-related tasks, including zero-shot classification, cross-modal retrieval, Visual Question Answering, and Multiple Choice Question. Our project is available at the Patho-R1 repository: https://github.com/Wenchuan-Zhang/Patho-R1.
@@ 1 Introduction
We present PathoCLIP, an open-source pathology adapted CLIP model, which outperforms state-of-the-art models in classification and retrieval tasks.
@@ 3.4 Reinforcement Learning
Reward Function Design. We design reward functions to guide model learning from structured and accurate responses. For the Format Reward, we require the model to format its response using <think>…</think> for intermediate reasoning steps and <answer>…</answer> for the final answer. A reward score of 1 is assigned if both tag pairs appear exactly once and no content exists outside these tags; otherwise, the score is 0. For the Accuracy Reward, the content within the <answer>…</answer> tags is extracted using regular expressions, and only exact matches to a valid option from the predefined set A, B, C, D, E, F are accepted. A reward score of 1 is assigned if the extracted answer matches the ground-truth label and 0 otherwise. For the Length-Aware Penalty, we implement a soft overlong punishment similar to the original work of DAPO. For more details on reward function design, see Appendix B.2 For GRPO, the reward function is:
@@ 4.4.1 Open-ended Benchmarks
Our proposed model, Patho-R1, is designed to perform step-by-step reasoning and produce accurate final answers. As shown in Figure 4, existing general-purpose and medical models do not necessarily possess the reasoning capability required to correctly answer pathology-specific questions. To evaluate its performance, we follow a fuzzy evaluation strategy in which Deepseek-R1 acts as the LLM-judge, scoring generated outputs based on ten well-defined criteria [41]. Evaluations are conducted using the VLMEvalKit [57] framework on two publicly available pathology VQA datasets: Quilt-VQA and Path-VQA. This setup ensures consistency, reproducibility, and fair comparison with other models. Our comprehensive evaluation highlights the strength of Patho-R1 in advancing multimodal reasoning in pathology. In addition to the traditional CoT prompts, we also used the latest Chain-of-Decision (CoD) prompts [58] to explore whether simplifying the reasoning chain affects model performance. For ablation studies and other experiments, see Appendix C.2 and C.3.
@@ 4.4.2 Close-ended Benchmarks
Closed-ended questions play a crucial role in pathology-related tasks, particularly in diagnostic classification. To evaluate model performance on such tasks, we consider two types of close-ended question datasets: (1) Yes/No questions, selected from PathVQA and QuiltVQA; and (2) multiple-choice questions, sourced from PathMMU [59], MedXpertQA [60], and OmniMedVQA [61]. For both types, we report accuracy as the evaluation metric to ensure consistency and comparability across datasets. Experimental results Table 3, Table 4 demonstrate that Patho-R1 outperforms previous state-of-the-art pathology-focused LMMs. Specifically, it surpasses PathGen-LLaVA-13B by 11.51 % on the PathMMU test-tiny split (PathGen-LLaVA-13B: 61.9%), by approximately 10% on the PathMMU test set, and by 5% on the PathMMU evaluation set.
@@ 5 Conclusion
In this study, we propose a comprehensive data curation pipeline for the three phases of pathology VLM training. Using this pipeline, we construct an extensive corpus containing 3.5 million image-text pairs, 500k SFT data with reasoning, and 10k MCQs for RL. Furthermore, we explore the end-to-end domain adaptation of pretrained VLMs, yielding Patho-R1 and PathoCLIP, two open-source models that contribute to the advancement of pathology-specific vision-language research and offer practical insights for future development in this domain.

### v3

@@ Abstract
Recent advances in vision-language models (VLMs) have enabled broad progress in the general medical field. However, pathology still remains a more challenging sub-domain, with current pathology-specific VLMs exhibiting limitations in both diagnostic accuracy and reasoning plausibility. Such shortcomings are largely attributable to the nature of current pathology datasets, which are primarily composed of image–description pairs that lack the depth and structured diagnostic paradigms employed by real-world pathologists. In this study, we leverage pathology textbooks and real-world pathology experts to construct high-quality, reasoning-oriented datasets. Building on this, we introduce Patho-R1, a multimodal RL-based pathology Reasoner, trained through a three-stage pipeline: (1) continued pretraining on 3.5 million image-text pairs for knowledge infusion; (2) supervised fine-tuning on 500k high-quality Chain-of-Thought samples for reasoning incentivizing; (3) reinforcement learning using Group Relative Policy Optimization and Decoupled Clip and Dynamic sAmpling Policy Optimization strategies for multimodal reasoning quality refinement. To further assess the alignment quality of our dataset, we propose Patho-CLIP, trained on the same figure-caption corpus used for continued pretraining. Comprehensive experimental results demonstrate that both Patho-CLIP and Patho-R1 achieve robust performance across a wide range of pathology-related tasks, including zero-shot classification, cross-modal retrieval, Visual Question Answering, and Multiple Choice Question. Our project is available at the Patho-R1 repository: https://github.com/Wenchuan-Zhang/Patho-R1.
@@ 1 Introduction
We present Patho-CLIP, an open-source pathology adapted CLIP model, which outperforms state-of-the-art models in classification and retrieval tasks.
@@ 3.4 Reinforcement Learning
Reward Function Design. We design reward functions to guide model learning from structured and accurate responses. For the Format Reward, we require the model to format its response using <think>…</think> for intermediate reasoning steps and <answer>…</answer> for the final answer. A reward score of 1 is assigned if both tag pairs appear exactly once and no content exists outside these tags; otherwise, the score is 0. For the Accuracy Reward, the content within the <answer>…</answer> tags is extracted using regular expressions, and only exact matches to a valid option from the predefined set A, B, C, D, E, F are accepted. A reward score of 1 is assigned if the extracted answer matches the ground-truth label and 0 otherwise. For the Length-Aware Penalty, we implement a soft overlong punishment similar to the original work of DAPO. For more details on reward function design, see Appendix B.2. For GRPO, the reward function is:
@@ 4.4.1 Open-ended Benchmarks
Our proposed model, Patho-R1, is designed to perform step-by-step reasoning and produce accurate final answers. As shown in Figure 4, existing general-purpose and medical models do not necessarily possess the reasoning capability required to correctly answer pathology-specific questions. To evaluate its performance, we follow a fuzzy evaluation strategy in which Deepseek-R1 acts as the LLM-judge, scoring generated outputs based on ten well-defined criteria [41]. Evaluations are conducted using the VLMEvalKit [57] framework on two publicly available pathology VQA datasets: Quilt-VQA and Path-VQA. This setup ensures consistency, reproducibility, and fair comparison with other models. Our comprehensive evaluation highlights the strength of Patho-R1 in advancing multimodal reasoning in pathology. In addition to the traditional CoT prompts, we also used the latest Chain-of-Draft (CoD) prompts [58] to explore whether simplifying the reasoning chain affects model performance.
@@ 4.4.2 Close-ended Benchmarks
Closed-ended questions play a crucial role in pathology-related tasks, particularly in diagnostic classification. To evaluate model performance on such tasks, we consider two types of close-ended question datasets: (1) Yes/No questions, selected from Path-VQA and Quilt-VQA; and (2) multiple-choice questions, sourced from PathMMU [59], MedXpertQA [60], and OmniMedVQA [61]. For both types, we report accuracy as the evaluation metric to ensure consistency and comparability across datasets. Experimental results Table 3, Table 4 demonstrate that Patho-R1 outperforms previous state-of-the-art pathology-focused LMMs. Specifically, it surpasses PathGen-LLaVA-13B by 7.63 % on the PathMMU test-tiny split (PathGen-LLaVA-13B: 61.9%), by approximately 5% on the PathMMU test set and PathMMU evaluation set. For ablation studies and other experiments, see Appendix C.2 and C.3.
@@ 5 Conclusion
In this study, we propose a comprehensive data curation pipeline for the three phases of pathology VLM training. Using this pipeline, we construct an extensive corpus containing 3.5 million image-text pairs, 500k SFT data with reasoning, and 10k MCQs for RL. Furthermore, we explore the end-to-end domain adaptation of pretrained VLMs, yielding Patho-R1 and Patho-CLIP, two open-source models that contribute to the advancement of pathology-specific vision-language research and offer practical insights for future development in this domain.